# Position: The Hidden Costs and Measurement Gaps of Reinforcement Learning with Verifiable Rewards

### v1

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) is a practical and scalable approach to enhancing large language models in areas such as math, code, and other structured tasks. Two questions motivate this paper: how much of the reported gains survive under strictly parity-controlled evaluation, and whether RLVR is cost-free or exacts a measurable tax. We argue that progress is real, but gains are often overstated due to three forces—an RLVR tax, evaluation pitfalls, and data contamination. Using a partial-prompt contamination audit and matched-budget reproductions across base and RL models, we show that several headline gaps shrink or vanish under clean, parity-controlled evaluation. We then propose a tax-aware training and evaluation protocol that co-optimizes accuracy, grounding, and calibrated abstention and standardizes budgeting and provenance checks. Applied to recent RLVR setups, this protocol yields more reliable estimates of reasoning gains and, in several cases, revises prior conclusions. Our position is constructive: RLVR is valuable and industry-ready; we advocate keeping its practical benefits while prioritizing reliability, safety, and measurement.
@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) has become a leading post-training route for improving large language models on math, code, and other structured tasks  (Luong et al., 2024; Wen et al., 2025a). By optimizing against automatically computable signals—unit tests for programs, exact numeric or string matches for math, or retrieval-grounded checks for citations—RLVR promises a scalable, label-efficient path to better reasoning. Recent results are striking: across multiple domains, RLVR systems often post large gains on standard benchmarks. Moreover, Figure 2 shows a rise in RLVR-tagged papers alongside improvements on AIME-24/25 through 2024–H1 2025, underscoring both the field’s momentum and the need to separate genuine reasoning gains from measurement and budgeting artifacts.
@@
Beneath the surface, however, sits a question that frames the current debate: does RLVR genuinely impart new reasoning capability, or does it mainly sharpen selection among behaviors the base model already knows how to produce? Parity-controlled studies show that base models can narrow or erase RLVR gaps when given matched sampling budgets—consistent with smarter search rather than capability expansion (Yue et al., 2025; Wu et al., 2025a). At the same time, several regimes report gains that are hard to recover by sampling alone: explicitly optimizing the multi-sample objective (pass@k training) (Wang et al., 2025b), curriculum via self-play with variational problem synthesis (Zhang et al., 2025a), and distribution-aware reward shaping that counters rank bias and diversity collapse (Liu et al., 2025b), alongside longer-horizon RL schedules and unlikeliness rewards (Liu et al., 2025a; He et al., 2025a). Even metric choice can flip conclusions: answer-only scores may diverge from process-aware metrics that require both a correct answer and a valid chain of thought.
@@
A second complication is what we call the RLVR tax: unintended costs that can ride along with apparent gains. In practice, RLVR tends to reduce abstention and increase stated confidence—sometimes even when answers are wrong—thereby shifting risk from “I don’t know” to assertive errors (Song et al., 2025; Mei et al., 2025; Yao et al., 2025). It can also chip away at instruction fidelity on longer generations, where adhering to formats or constraints becomes harder as chains grow (Fu et al., 2025; Li et al., 2025). Finally, longer, more explicit reasoning traces expand the attack and leakage surface, raising jailbreak success and privacy exposure if left unchecked (Zhou et al., 2025; Jiang et al., 2025; Ackerman & Panickssery, 2025; Green et al., 2025; Huang et al., 2025; Zhang et al., 2025b).
@@
A third, orthogonal complication is measurement. Reported advances are sensitive to sampling budgets and decoding settings, to the stability of LLM-as-a-judge evaluators, to calibration drift, and to data provenance (Hochlehnert et al., 2025; Chandak et al., 2025; Brown et al., 2024; Muennighoff et al., 2025; Zhao et al., 2025b; Wang et al., 2024a; b; Leng et al., 2024; Shen et al., 2025; Wu et al., 2025b; He et al., 2025b; Mirzadeh et al., 2024; Liu et al., 2024b). When budgets are matched, judges are stress-tested, and datasets are versioned and decontaminated, reported improvement gaps can diminish, suggesting that part of the apparent progress reflects evaluation design rather than durable capability (Hochlehnert et al., 2025; Chandak et al., 2025). To help orient the reader, Figure 1 summarizes three threads (taxes, evaluation pitfalls, and contamination) and previews the tax-aware protocol we adopt.
@@
Throughout, we use RLVR to mean post-training that optimizes LLMs against automatically checkable signals rather than human preference models (Ouyang et al., 2022). Typical implementations adopt PPO/GRPO/DAPO-style updates with a KL penalty to the base policy, often mixing offline rollouts with online sampling, adding entropy regularization, and using selective filtering or gating. Rewards are frequently componentized—for correctness, grounding/citation sufficiency, and calibrated refusal—and introduced in stages. Our analysis centers on open-weight models fine-tuned for math/code/question answering (QA) with verifiable objectives, in both single- and multi-domain regimes. We deliberately avoid a catalogue of low-level algorithmic variants except where they bear on reliability (e.g., unlikeliness rewards that counter rank bias, or curricula that change variance). The primary threats considered are privacy leakage from long chains of thought (CoT) and increased jailbreak susceptibility during or after RLVR.
Figure 2: Monthly RLVR activity vs. AIME performance (time span: May 2024–June 2025). Left axis: count of pages per month whose title or abstract contains “RLVR” or “reinforcement learning with verifiable rewards” (Google Scholar and arXiv via SerpAPI). Right axis: pass@1 (%) on AIME-24 (circles, warm) and AIME-25 (squares, cool) for selected models; labels show model names.
Position and Roadmap.
RLVR is valuable and industry-ready, but we encourage prioritizing reliability, safety, and measurement. We (i) synthesize evidence on sharpening vs. expansion (section 2); (ii) analyze the RLVR tax and how it distorts reported improvements (section 3); (iii) reproduce representative gaps under parity controls (section 4); (iv) audit contamination with partial-prompt probes (section 5); and (v) offer a tax-aware protocol that co-optimizes accuracy, grounding, and calibrated abstention while standardizing budgeting and provenance checks (section 6).
@@ 2 What Changes Under RLVR? An Evidence Overview
We review empirical signals about RLVR’s effects without taking a side on the “expansion v.s. sharpening” question. Skeptical and optimistic perspectives recur in the literature, and each tends to appear under identifiable conditions.
@@ 2.1 Skeptical Lens: Sharpening and Sampling Effects
Under a skeptical lens, RLVR often appears to improve sample efficiency and steer the model toward high-reward regions already present in the base distribution, rather than broaden fundamental capability (Gandhi et al., 2025; Shah et al., 2025). When base models are evaluated under matched sampling budgets (i.e., pass@k), gaps between base and RLVR-trained models often shrink (Yue et al., 2025), consistent with selection rather than acquisition. Empirically, outputs can collapse toward a dominant pretraining mode with reduced diversity (Zhao et al., 2025a), and theory suggests that after RLVR, support shrinkage tends to outweigh support expansion (Wu et al., 2025a). Stress tests add caution: performance collapses at higher complexities have been reported even with ample compute (Shojaee et al., 2025), although some failures trace to evaluation artifacts such as token limits and unsolvable items, underscoring the need for budget parity, seed control, and dataset hygiene (see Section 4(Opus & Lawsen, 2025). Finally, several works separate knowledge injection from policy optimization: pipelines that use SFT or distillation to inject knowledge and RLVR to select among candidates typically beat pure GRPO, suggesting RLVR optimizes within learned support (Ma et al., 2025; Wan et al., 2025; Chen et al., 2025a); guidance and self-distillation further help over vanilla GRPO (Nath et al., 2025; Liu et al., 2025c).
@@ 2.2 Optimistic Lens: Signals of Capability Broadening
Under an optimistic lens, carefully designed RLVR appears to expand what models can do, not merely sharpen selection. Prolonged optimization with explicit KL control and periodic resets (ProRL) reports trajectories that are not reachable by sampling the base policy alone, with the largest gains where the base initially struggles (Liu et al., 2025a). Reward shaping also matters: by diagnosing a GRPO tendency to reinforce high-probability completions, He et al. (2025a) show that adding unlikeliness rewards and training deeper widens the model’s support and improves multi-sample theorem proving, indicating qualitative changes to reachable reasoning modes. These effects are conditioned by pretraining coverage: cross-domain analyses find broader benefits in Math/Code/Science, while under-represented areas such as Logic/Simulation/Tabular realize meaningful gains when RLVR is applied in-domain (Cheng et al., 2025b). Consistent with this, process-aware metrics that require both a correct answer and a syntactically valid chain (CoT-pass@k) often reveal larger RLVR deltas than answer-only measures (Wen et al., 2025b). While generalization to substantially harder regimes remains an open goal (Sun et al., 2025b), the weight of evidence supports a constructive view: with sufficient optimization depth, bias-aware rewards, and the right domain coverage, RLVR can broaden reasoning capabilities.
@@
Answer-only metrics (e.g., pass@k) can show muted gains, whereas process-aware criteria such as CoT-pass@k—which require both a correct answer and a syntactically valid chain—often reveal larger deltas in settings explicitly optimized for process rewards (Wen et al., 2025b). We treat these as behavioral signals rather than explanations: chains need not be faithful to internal computation (Chen et al., 2025d). Accordingly, we report both answer- and process-aware metrics alongside calibration (ECE/abstention), and we prefer budget-matched saturation curves over single-point pass@k (Section 4).
@@
Across both perspectives, two themes are consistent: (i) the base model’s strength and coverage strongly condition observed gains, and (ii) reward/metric design and evaluation protocol (sampling budgets, seeds, and contamination hygiene) can change the sign of conclusions. These observations motivate our stance: treat reliability, safety, and measurement as first-class objectives. The following sections detail the RLVR tax (Section 3), show how evaluation choices inflate or erode claims (Section 4), and present a unified, tax-aware training and evaluation protocol that yields more trustworthy estimates and more transferable improvements (Section 6).
3 Perspective I — The RLVR Tax: Benefits, Costs, and Controls
RLVR reliably lifts accuracy on verifiable tasks, but those gains are easy to misread if we ignore systematic costs—the RLVR tax. We focus on three recurrent pressures: (i) hallucination and overconfidence, (ii) erosion of instruction following, and (iii) safety and privacy exposure. Each maps directly to our broader thesis: without tax-aware training and sober measurement, reported “reasoning gains” are easy to overstate and hard to transfer.
@@ 3.1 Hallucination and Overconfidence
RLVR can suppress refusals while amplifying confident errors. Empirically, refusal rates often collapse after RLVR, shifting abstentions into assertive answers (Song et al., 2025); models may also repeat flawed steps or produce CoT that diverges from the final answer even when accuracy rises (Yao et al., 2025). A mechanistic reading is that sparse, verifiable rewards combined with entropy pressure encourage determinacy under weak evidence, producing high-variance gradients and spurious local optima (Li & Ng, 2025). Consistent with this, multiple studies find substantial miscalibration: self-reported confidence remains high even on incorrect responses and can increase with longer CoT traces (Mei et al., 2025; Zeng et al., 2025; Kirichenko et al., 2025).
We make this concrete with a factual-QA control that tracks Not attempted (lower is better), Accuracy (shared)—accuracy on items that both models at a given scale attempted (higher is better)—and Expected Calibration Error (ECE; lower is better) computed on shared items from the model’s stated confidence (10 uniform bins; see section 6 (C) for details). The dominant effect is attempt inflation: stronger RLVR policies attempt far more questions. For example, moving from DeepSeek-V3 to DeepSeek-R1 (Guo et al., 2025) sharply reduces abstentions and improves shared ECE, but accuracy on the newly attempted tail is modest, so reported gains depend on how one aggregates across attempted vs. unattempted items (Zeng et al., 2025). In other words, scaling RL encourages attempts; whether this yields net quality improvements depends on calibration and gating.
In practice, we mitigate these effects by (i) decomposing reward into correctness, grounding/citation sufficiency, and calibrated refusal (see section 6 (A)); (ii) adding calibration gates (ECE/entropy) and early stopping when confidence outpaces accuracy (see section 6 (C)); and (iii) reporting ECE and refusal alongside accuracy, plus shared-accuracy and attempt-aware summaries (see section 6 (D)).
@@ 3.2 Instruction Following
Optimizing hard for reasoning can erode controllability—especially for long generations. Several studies observe regressions in instruction fidelity when training emphasizes extended chain-of-thought or purely verifiable endpoints (Fu et al., 2025; Li et al., 2025). Moreover, recent evidence shows that models tend to overfit to a small set of verifiable constraints and struggle to generalize; IFBench (58 out-of-domain constraints) documents this gap and finds that RLVR improves precise instruction-following generalization (Pyatkin et al., 2025). We counter this by making format/constraint adherence a first-class term in the reward (see section 6 (A)) and by alternating objectives or schedules so the model periodically re-anchors to instruction faithfulness rather than overfitting to long-CoT trajectories (see section 6 (B)).

### v2

# The Hidden Costs and Measurement Gaps of Reinforcement Learning with Verifiable Rewards

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) is a practical, scalable way to improve large language models on math, code, and other structured tasks. However, we argue that many headline RLVR gains are not yet well validated because reports often conflate policy improvement with three confounds: (i) budget mismatch between RLVR and baseline evaluation, (ii) attempt inflation and calibration drift that convert abstentions into confident answers, and (iii) data contamination in benchmarks. Using budget-matched reproductions and partial-prompt contamination probes, we find that several widely cited gaps shrink substantially or disappear once budgets, prompts, and dataset versions are matched, and contaminated sets are treated as memorization probes rather than evidence of reasoning. This does not mean that RLVR is ineffective, but it implies that current measurements often overstate capability gains and obscure reliability costs. We therefore propose a compact, tax-aware minimum standard for RLVR training and evaluation: budget-matched saturation curves with variance, calibration, and abstention tracking, one judge robustness stress test when LLM judges are used, and an explicit contamination screen. With these controls, RLVR remains effective and deployable in verifiable domains, but reasoning gains should be treated as provisional without them.
The Hidden Costs and Measurement Gaps of Reinforcement Learning with Verifiable Rewards
Fang Wu1,*†   Aaron Tu2,*   Weihao Xuan3,*   Heli Qi4,*   Xu Huang5   Qingcheng Zeng6 Shayan Talaei1   Yijia Xiao7   Peng Xia8   Xiangru Tang9   Yuchen Zhuang5   Bing Hu10 Hanqun Cao11   Wenqi Shi12   Rui Yang14   Nan Liu14   Huaxiu Yao8   Ge Liu13 Li Erran Li15   Amin Saberi1   Naoto Yokoya3   Jure Leskovec1   Yejin Choi1,† 1Stanford University  2UC Berkeley  3The University of Tokyo  4Waseda University 5Georgia Tech  6Northwestern University  7UCLA  8UNC Chapel Hill  9Yale University 10Independent Researcher  11CUHK  12UT Southwestern Medical Center  13UIUC 14National University of Singapore  15Amazon AWS AI *Equal Contribution   Corresponding Authors
@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) has become a leading post-training route for improving large language models on math and code tasks (Luong et al., 2024; Wen et al., 2025a). By optimizing against automatically computable signals such as unit tests for programs, exact numeric or string matches for math, or retrieval-grounded checks for citations, RLVR promises a path to better performance. Recent work reports large benchmark gains across domains. Appendix Figure 2 shows a trend plot of RLVR-tagged paper activity and AIME-24/25 performance with striking monthly gains. Our position is that RLVR is effective and deployable for verifiable domains, but “reasoning” gains should be treated as provisional without a small core of tax-aware controls.
@@
A central question that frames the current debate is: Does RLVR genuinely impart new reasoning capability, or does it mainly sharpen selection among behaviors the base model already knows how to produce? Budget-parity controlled studies show that base models can narrow or erase RLVR gaps when given matched sampling budgets, consistent with smarter search rather than capability expansion (Yue et al., 2025; Wu et al., 2025c). At the same time, several regimes report gains that are hard to recover by sampling alone: explicitly optimizing the multi-sample objective (pass@k training) (Wang et al., 2025), curriculum via self-play with variational problem synthesis (Zhang et al., 2025a), and distribution-aware reward shaping that counters rank bias and diversity collapse (Liu et al., 2025b), alongside longer-horizon RL schedules and unlikeliness rewards (Liu et al., 2025a; He et al., 2025a). Metric choice can flip conclusions: answer-only scores may diverge from process-aware metrics that require both a correct answer and a valid chain of thought.
@@
A second complication is what we call the RLVR tax: unintended empirical side effects that ride along with apparent gains under current reasoning-style post-training (reduced abstention, miscalibration, instruction-fidelity drift, and a larger safety/privacy surface due to longer traces). This tax is not inevitable and not unique to RLVR; similar patterns arise under reasoning-heavy SFT and RLHF (Wu et al., 2025b). We focus on RLVR because verifiable objectives and open weights make these trade-offs measurable at relatively low cost. In practice, RLVR tends to reduce abstention and increase stated confidence, sometimes even when answers are wrong, thereby shifting risk from “I don’t know” to assertive errors (Song et al., 2025; Mei et al., 2025; Yao et al., 2025). It can also chip away at instruction fidelity on longer generations, where adhering to formats or constraints becomes harder as chains grow (Fu et al., 2025; Li et al., 2025). Finally, longer, more explicit reasoning traces expand the attack and leakage surface, raising jailbreak success and privacy exposure if left unchecked (Zhou et al., 2025a; Jiang et al., 2025; Ackerman and Panickssery, 2025; Green et al., 2025; Huang et al., 2025; Zhang et al., 2025b).
@@
A third complication is measurement. Reported advances are sensitive to sampling budgets and decoding settings, to the stability of LLM-as-a-judge evaluators, to calibration drift, and to data provenance (Hochlehnert et al., 2025; Chandak et al., 2025; Brown and others, 2024; Muennighoff et al., 2025; Zhao et al., 2025; Wang et al., 2024a, b; Leng et al., 2024; Shen et al., 2025; Ye et al., 2026; He et al., 2025b; Mirzadeh et al., 2024; Liu et al., 2024b; Xu et al., 2025; Wu and others, 2024; Wu et al., 2025a, ). When budgets are matched, judges are stress-tested, and datasets are versioned and decontaminated, reported improvement gaps can diminish, suggesting that part of the apparent progress reflects evaluation design rather than durable capability (Hochlehnert et al., 2025; Chandak et al., 2025). We organize the paper around three threads: the RLVR tax, evaluation pitfalls, and data contamination, culminating in a compact tax-aware minimum standard. A visual roadmap is provided in Figure 1.
@@
We use RLVR to mean post-training that optimizes LLMs against automatically checkable signals rather than human preference models (Ouyang and others, 2022). Typical implementations adopt PPO/GRPO/DAPO-style updates with a KL penalty to the base policy, often mixing offline rollouts with online sampling, adding entropy regularization, and using selective filtering or gating. Rewards are frequently componentized (for correctness, grounding/citation sufficiency, and calibrated refusal) and introduced in stages. Our analysis centers on open-weight models fine-tuned for math/code/question answering (QA) with verifiable objectives, in both single- and multi-domain regimes. We deliberately avoid a catalogue of low-level algorithmic variants except where they bear on reliability (e.g., unlikeliness rewards that counter rank bias, or curricula that change variance). The primary threats considered are privacy leakage from long chains of thought (CoT) and increased jailbreak susceptibility during or after RLVR.
Contributions and Roadmap.
Our contributions are threefold: (1) Measurement: we provide budget-matched reproductions and a unified gap table that isolates from budget mismatch, judge and template drift, and dataset versioning (Table 2). (2) Tax: we quantify attempt inflation and the resulting calibration and refusal costs under RLVR and matched SFT controls (Table 1). (3) Protocol: based on these findings, we distill a compact tax-aware minimum standard for RLVR training and evaluation, intended as a minimum credible reporting bar, instantiated in section 6. The paper is organized accordingly: we (i) synthesize evidence on sharpening vs. expansion (section 2); (ii) analyze the RLVR tax and how it distorts reported improvements (section 3); (iii) reproduce representative gaps under parity controls (section 4); (iv) audit contamination with partial-prompt probes (section 5); and (v) present the tax-aware minimum standard (section 6).
@@ 2 What Changes Under RLVR? An Evidence Overview
We summarize evidence under two lenses: sharpening vs broadening. Skeptical and optimistic perspectives recur in the literature, and each tends to appear under identifiable conditions.
@@ 2.1 Skeptical Lens: Sharpening and Sampling Effects
RLVR often appears to improve sample efficiency and steer the model toward high-reward regions already present in the base distribution, rather than broaden fundamental capability (Gandhi et al., 2025; Shah and others, 2025). When base models are evaluated under matched sampling budgets (i.e., pass@k), gaps between base and RLVR-trained models often shrink (Yue et al., 2025), consistent with sharpening rather than expansion. Empirically, outputs can collapse toward a dominant pretraining mode with reduced diversity (Zhao and others, 2025), and theory suggests that after RLVR, support shrinkage tends to outweigh support expansion (Wu et al., 2025c). Stress tests add caution: performance collapses at higher complexities have been reported even with ample compute (Shojaee et al., 2025), although some failures trace to evaluation artifacts such as token limits and unsolvable items, underscoring the need for budget parity, seed control, and dataset hygiene (see section 4(Opus and Lawsen, 2025). Finally, several works separate knowledge injection from policy optimization: pipelines that use SFT or distillation to inject knowledge and RLVR to select among candidates typically beat pure GRPO, suggesting RLVR optimizes within learned support (Ma et al., 2025; Wan et al., 2025; Chen et al., 2025a); guidance and self-distillation further help over vanilla GRPO (Nath et al., 2025; Liu et al., 2025c).
@@ 2.2 Optimistic Lens: Signals of Capability Broadening
Carefully designed RLVR appears to expand what models can do, not merely sharpen selection. Prolonged optimization with explicit KL control and periodic resets (ProRL) reports trajectories that are not reachable by sampling the base policy alone, with the largest gains where the base initially struggles (Liu et al., 2025a). Reward shaping also matters: by diagnosing a GRPO tendency to reinforce high-probability completions, He et al. (2025a) shows that adding unlikeliness rewards and training deeper widens the model’s support and improves multi-sample theorem proving, indicating qualitative changes to reachable reasoning modes. These effects are conditioned by pretraining coverage: cross-domain analyses find broader benefits in Math/Code/Science, while under-represented areas such as Logic/Simulation/Tabular realize meaningful gains when RLVR is applied in-domain (Cheng et al., 2025). Consistent with this, process-aware metrics that require both a correct answer and a syntactically valid chain (CoT-pass@k) often reveal larger RLVR deltas than answer-only measures (Wen et al., 2025b). While generalization to substantially harder regimes remains an open goal (Sun et al., 2025b), the weight of evidence supports a constructive view: with sufficient optimization depth, bias-aware rewards, and the right domain coverage, RLVR can broaden reasoning capabilities.
@@
Answer-only metrics (e.g., pass@k) can show muted gains, whereas process-aware criteria such as CoT-pass@k (which require both a correct answer and a syntactically valid chain) often reveal larger deltas in settings explicitly optimized for process rewards (Wen et al., 2025b). We treat these as behavioral signals rather than explanations: chains need not be faithful to internal computation (Chen et al., 2025c). Accordingly, we report both answer- and process-aware metrics alongside calibration (ECE/abstention), and we prefer budget-matched saturation curves over single-point pass@k (section 4).
@@
Across both perspectives, two themes are consistent: (i) the base model’s strength and coverage strongly condition observed gains, and (ii) reward/metric design and evaluation protocol (sampling budgets, seeds, and contamination hygiene) can change the sign of conclusions. These observations motivate our position.
3 Perspective I: The RLVR Tax: Benefits, Costs, and Controls
RLVR reliably lifts accuracy on verifiable tasks, but those gains are easy to misread if we ignore systematic costs, the RLVR tax. We elaborate on three recurrent pressures: (i) hallucination and overconfidence, (ii) erosion of instruction following, and (iii) safety and privacy exposure. Each maps directly to our position: without tax-aware controls and sober measurement, reported “reasoning gains” are easy to overstate and hard to transfer.
@@ 3.1 Hallucination and Overconfidence
RLVR can suppress refusals while amplifying confident errors. Empirically, refusal rates often collapse after RLVR, shifting abstentions into assertive answers (Song et al., 2025); models may also repeat flawed steps or produce CoT that diverges from the final answer even when accuracy rises (Yao et al., 2025). One hypothesis is that sparse, verifiable rewards combined with entropy pressure tend to encourage determinacy under weak evidence, producing high-variance gradients and spurious local optima (Li and Ng, 2025). Consistent with this, multiple studies find substantial miscalibration: self-reported confidence remains high even on incorrect responses and can increase with longer CoT traces (Mei et al., 2025; Zeng et al., 2025; Kirichenko et al., 2025).
We make this concrete with a factual-QA control that tracks Not attempted (lower is better), Accuracy (shared) (accuracy on items that both models at a given scale attempted, higher is better), and Expected Calibration Error (ECE; lower is better) computed on shared items from the model’s stated confidence. The dominant effect is attempt inflation: stronger RLVR policies attempt far more questions. For example, moving from DeepSeek-V3 to DeepSeek-R1 (Guo et al., 2025) sharply reduces abstentions and improves shared ECE, but accuracy on the newly attempted tail is modest, so reported gains depend on how one aggregates across attempted vs. unattempted items (Zeng et al., 2025). In other words, scaling RL encourages attempts; whether this yields net quality improvements depends on calibration and gating.
@@ 3.2 Instruction Following
Reasoning-optimized training can erode controllability, especially for long generations. Several studies observe regressions in instruction fidelity when training emphasizes extended chain-of-thought or purely verifiable endpoints (Fu et al., 2025; Li et al., 2025). Moreover, recent evidence shows that models tend to overfit to a small set of verifiable constraints and struggle to generalize; IFBench (58 out-of-domain constraints) documents this gap and finds that RLVR improves precise instruction-following generalization (Pyatkin et al., 2025). This is why our minimum standard in section 6 assumes either an explicit instruction or format component in the reward, or at least a separate instruction-following evaluation pack.